Show that the density of the nucleus is 2.3×1017 kgm3
1
Expert's answer
2020-06-12T11:16:05-0400

The radius of the nucleus can be expressed in terms of mass number:


"r=r_0A^{1\/3}."

The volume of the nucleus:


"V=\frac{4}{3}\pi r^3=\frac{4}{3}\pi r_0^3A."

The mass in kilograms in terms of A:

"m=A\cdot1.66\cdot10^{-27}."

The nuclear density


"\rho=\frac{m}{V}=\frac{3A\cdot1.66\cdot10^{-27}\text{ kg}}{4\pi r_0^3A}=\frac{3\cdot1.66\cdot10^{-27}\text{ kg}}{4\pi r_0^3},\\\space\\
\rho=\frac{3\cdot1.66\cdot10^{-27}\text{ kg}}{4\cdot3.14 (1.2\cdot10^{-15})^3}=2.29\cdot10^{17}\text{ kg/m}^3."